# Is Less More? Exploring Token Condensation as Training-free Adaptation for CLIP

### v2

@@ 5 Conclusion
In this paper, we introduced Token Condensation as Adaptation (TCA), a novel training-free test-time adaptation method for CLIP models. Our comprehensive experiments demonstrated that token condensation significantly benefits the visual-text alignment in CLIP, which can further serve as an interpretation of visual semantics. Additionally, our method reduces GFLOPs as a beneficial byproduct, enhancing computational efficiency. For fair comparisons, we fixed the GFLOPs budget by pre-setting $R$ in our experiments; however, the condensing rate can be adaptively estimated using the distances to the merging center as an indicator, which can be a promising direction for future research. We also acknowledge the limitations of TCA as a training-free method, particularly its bottleneck in handling datasets with severe distribution shifts (refer to the supplementary material).
@@
This supplementary material provides additional descriptions of the proposed TCA, including empirical results and algorithm. Visual aids for token condensation are also included to enhance understanding of the proposed method.
@@ 5.1 Additional Results
Comparison with State-of-the-Art TTA Using Token Condensation. We additionally evaluate the performance when combining TDA with token pruning and merging baselines and show the performance gain over TDA $+$ ATS in Fig. 6. Although TDA achieves considerable performance gain, it heavily relies on the negative cache and a large set of hyperparameters. In contrast, TCA’s accuracy gain significantly surpasses that of TDA $+$ EViT and TDA $+$ ToME across multiple datasets and on average, even with a minimal set of hyperparameters, highlighting its superior adaptation capability.
@@ 5.2 Algorithm
7:Domain <cls> Token Selection: Obtain $\mathbf{A}^{l-1}_{c^{*}}$ by Eq. 8, using domain <cls> tokens in $\mathbf{\mathfrak{R}}$ and sample’s <cls> token $\mathbf{v}^{l}_{\text{cls}}$
@@
10:       Token Pruning: Obtain $\mathbf{\hat{V}}^{l}_{\text{prune}}$ via LABEL:eq:sort
@@
Algorithm 1 outlines the process for performing token pruning and merging at layer $l$ in a ViT-based CLIP model. We first obtain the averaged domain <cls> tokens $\mathbf{A}^{l-1}_{c^{*}}$ by the <cls> tokens saved in the reservoir $\mathfrak{R}$. Token condensation is then conducted given the domain <cls> token. Specifically, we conduct token pruning by relative ranking positions of token $i$ across multiple attention heads. Then, coreset selection is used for token merging. Finally, we concatenate the <cls> token $\mathbf{v}^{l}_{\operatorname{cls}}$ with the retained tokens as the input for the next layer, where the original $N+1$ tokens are shrunk to $(R\cdot N)+1$, thereby reducing the computational cost.

### v3

@@ 5 Conclusion
We introduced Token Condensation as Adaptation (TCA), a novel training-free test-time adaptation method on VLMs such as CLIP. Our experiments across various VLMs demonstrated that token condensation significantly enhances visual-text alignment while also serving as an interpretation of visual semantics. Besides, TCA reduces GFLOPs as a beneficial byproduct, improving computational efficiency. Despite our empirical findings, a rigorous theory of CLIP’s generalizability remains underexplored. To bridge the gap, we provide a theoretical analysis, examine TCA’s applicability to other VLMs, and acknowledge its limitations in the supplementary material.
@@
This supplementary material provides additional details of TCA, including method descriptions, theoretical analysis, empirical results, and the algorithm. We also discuss TCA’s applicability and limitations. To further illustrate the method, we include visual aids for token condensation.
  • Sec. 5.1: Details of the Coreset Selection Strategy.
  • Sec. 5.2: Theoretical Analysis.
@@
5.1 Details of Coreset Selection
In domain-aware token merging, we first identify the most representative tokens $\hat{\mathbf{V}}^{l}_{\operatorname{merge}}\in\mathbb{R}^{K\times D_{v}}$ from $\mathbf{V}^{l}_{\Phi}$ and assigns the remaining ambiguous tokens to these selected tokens. This strategy is equivalent to solving the K-Center problem [62, 46]. The objective is to select $K$ center tokens such that the maximum distance between any token and its nearest center is minimized. The greedy search for coreset optimization is defined as follows:
where $\mathcal{N}(k)$ represents the set of tokens assigned to center $k$. The value of $K$ is kept small, with $K\ll N$, allowing our merging algorithm to operate with linear complexity.
5.2 Theoretical Analysis
The theoretical foundations of CLIP’s generalization remain underexplored, with ongoing debates on whether it arises from train-test similarity [34], spurious feature reliance [56], or other factors. While we did not include rigorous proof, we connect our TCA to PAC-Bayesian generalization theory. We model token selection as a stochastic hypothesis, where the posterior $\mathbb{Q}$ over retained tokens follows a Gibbs formulation, favoring subsets that minimize cosine similarity variance with texts:
$\footnotesize\begin{split}\mathbb{Q}(\hat{\mathbf{V}})&=\frac{1}{Z}\exp\left(-\lambda\mathrm{Var}\left(\operatorname{cos}(\mathbf{V},\mathbf{t}_{c})\right)\right),\\ \mathbb{E}_{\operatorname{ood}}[-\operatorname{cos}(\hat{\mathbf{V}},\mathbf{t}_{c})]\leq&\mathbb{E}_{\operatorname{id}}[-\operatorname{cos}(\mathbf{V},\mathbf{t}_{c})]+\sqrt{\frac{1}{2}(D_{\operatorname{KL}}(\mathbb{Q}\|\mathbb{P})+\log\frac{m}{\delta}}.\end{split}$
This supports the PAC-Bayes bound, where TCA improves generalization by reducing KL divergence between test-time token selection and CLIP’s inaccessible pretraining distribution, which we approximate using DTR. Empirical results in Fig. 2(b) confirm this, showing that retained tokens act as stable anchors for text alignment.
@@ 5.3 Additional Results
Table 7: Impact of scale factor $\beta$.
$\beta$ 0.01 0.05 1 3 5
Pets 89.51 89.53 89.37 89.42 89.26
Flower102 73.33 73.08 70.93 70.56 70.44
EuroSAT 63.64 64.06 69.86 70.26 70.43
Impact of Logits Correction Temperature $\beta$. Tab. 7 examines how different logits correction temperatures $\beta$ affect the adaptation results. The intuition is that with a smaller $\beta$ value, the logits correction will emphasize the tokens in shallower layers (Eq. 9), while a larger $\beta$ value will shift the focus to deeper layers. We observe that a smaller value of $\beta$ is preferred for the Pets dataset as it contains animals as objects, requiring more high-level contextual information for accurate predictions [40]. In contrast, for EuroSAT, the best predictions are obtained with larger $\beta$ values, suggesting that low-level, local information is crucial. This aligns well with the nature of the dataset, where different types of land can be distinguished by features such as colors and edges. Nevertheless, our method consistently provides significant improvements across all $\beta$ values, with accuracy gains of up to 20%, highlighting the effectiveness of logits correction using the domain anchor tokens.
Table 8: Impact of correction weight $\lambda$.
$\lambda$ 2 3 4 5 6 7 8
Pets 89.53 89.32 89.13 88.96 88.96 88.66 88.44
Flower102 72.43 72.76 73.20 73.16 73.16 73.33 73.16
EuroSAT 60.15 65.74 68.80 69.51 69.84 70.16 70.43
@@ 5.4 Algorithm
7:Domain Anchor Token Selection: Obtain $\mathbf{A}^{l-1}_{c^{*}}$, using domain anchor tokens in $\mathbf{\mathfrak{R}}$ and sample’s <cls> token $\mathbf{v}^{l}_{\text{cls}}$
@@
10:       Token Pruning: Obtain $\mathbf{\hat{V}}^{l}_{\text{prune}}$ via Eq. 7
@@
Algorithm 1 outlines the process for performing token pruning and merging at layer $l$ in a ViT-based CLIP model. We first obtain the averaged domain anchor tokens $\mathbf{A}^{l-1}_{c^{*}}$ by the <cls> tokens saved in the reservoir $\mathfrak{R}$. Token condensation is then conducted given the domain anchor token. Specifically, we conduct token pruning by relative ranking positions of token $i$ across multiple attention heads. Then, coreset selection is used for token merging. Finally, we concatenate the <cls> token $\mathbf{v}^{l}_{\operatorname{cls}}$ with the retained tokens as the input for the next layer, where the original $N+1$ tokens are shrunk to $(R\cdot N)+1$, thereby reducing the computational cost.
@@ 5.5 Quantitative Study
5.6 Discussion on TCA’s Generalizability
TCA is designed for VLMs such as CLIP, SigLIP, and SigLIP v2, requiring only minor modifications. These models share a key characteristic: they compute cosine similarity between modalities for zero-shot image classification. For CLIP, we use the <cls> token as a guiding indicator throughout the method. In contrast, for the SigLIP series, we take the average over attention weights since their architecture does not include a visual <cls> token. The way we determine the domain anchor token and perform token condensation is inherently tied to how each VLM extracts visual features for alignment. We acknowledge that TCA may not directly apply to models like LLaVA [28], as they are not designed for cross-modal alignment but rather for text generation, dictated by their architectural constraints. While this limits direct applicability, it does not diminish TCA’s effectiveness in its intended scope. Adapting it to such models would likely require a fundamental architectural redesign.